Simplify the following expression.

sqrt-104 sqrt-8

Answer by stanbon(75887)

sqrt-104 sqrt-8
----
Assuming you mean to multiply you get:
sqrt(104)(sqrt(-1))(sqrt(8))(sqrt(-1)
---------------------
= sqrt(4*26)i * sqrt(4*2)i
---
= i^2* 2sqrt(26)*2*sqrt(2)
---
= -4sqrt(52)
----
= -4*2*sqrt(13)
=====================
Cheers,
Stan H.